OpenStudy (anonymous):

find the domain: f(y)= y + 6 / y^2 -6y

OpenStudy (anonymous):

is just y^2 on the bottom or the entire statement (y^2 -6y) on the bottom?

OpenStudy (anonymous):

just bottom.

OpenStudy (anonymous):

\[y + 6 \div y ^{2}-6y\]

OpenStudy (hba):

\[f(y)=\frac{ y+6 }{ y^2-6y }\] Or \[f(y)=\frac{ y+6 }{ y^2 }-6y\]

OpenStudy (anonymous):

interval notation as answer

OpenStudy (anonymous):

the first one!

OpenStudy (anonymous):

with that at the bottom factor out a y out of the bottom to get y(y-6) now the equation does not exist when the denominator is zero (either y or y-6 is zero) Use this to build to equations y=0 and y-6=0 to get your answers y=0 y=6 These are the points where the equation is undefined so the interval would be (-infinity,0) U (0,6) U (6,infinity) the round brackets are very important as they mean that the equation does not exist at precisely those numbers

OpenStudy (anonymous):

okay so they are all round brackets... when do I use these? [ ]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

if the equation exists at the number closest to the bracket then use a square bracket if it doesn't then use a round bracket. For positive and negative infinity always use a round bracket